Harvey has some 1 dollar bills and some 5 dollar bills. In all he has a total of 6 bills worth 22 dollars. Let x represent the number of 1 dollar bills and y represent the number of 5 dollar bills. Write a system of equations the information and use to determine how many of each denomination Harvey has.
Answer by stanbon(75887)

In all he has a total of 6 bills worth 22 dollars. Let x represent the number of 1 dollar bills and y represent the number of 5 dollar bills. Write a system of equations the information and use to determine how many of each denomination Harvey has.
------------------------
Quantity Equation: x + y = 6
Value Equation...: x + 5y= 22
-----------------------------------
Subtract the 1st from the 2nd to get:
4y = 16
y = 4 (# of $5 bills)
Since x+y = 6, x = 2 (# of $1 bills)
Cheers,
Stan H.